# Aligning Cyber Space with Physical World: A Comprehensive Survey on Embodied AI

### v7

@@ IV-B1 Two-stage 3D Visual Grounding methods
Existing 3D VG methods often rely on extensive labeled data for training or show limitations in processing complex language queries. Inspired by the impressive language understanding capabilities of LLMs, LLM-Grounder [138] proposed an open vocabulary 3D visual grounding pipeline that requires no labeled data, leveraging LLM to decompose queries and generate plans for object identification, followed by evaluating spatial and commonsense relations to select the best matching object. To capture view-dependent queries and decipher spatial relations in 3D space, ZSVG3D [139] designed a zero-shot open-vocabulary 3D visual grounding method that uses LLM to identify relevant objects and perform reasoning, transforming this process into a scripted visual program and then into executable Python code to predict object locations.
However, as shown in Fig. 8 (b), these two-stage methods face the dilemma of determining the number of proposals because the 3D detectors in the first stage require sampling keypoints to represent the entire 3D scene and generate corresponding proposals for each keypoint. Sparse proposals may overlook targets in the first stage, making them unmatchable in the second stage. Conversely, dense proposals may contain inevitable redundant objects, leading to difficulties in distinguishing targets in the second stage due to overly complex inter-proposal relationships. Moreover, the keypoint sampling strategy is language-agnostic, which increases the difficulty for detectors to identify language-related proposals.
IV-B2 One-stage 3D Visual Grounding methods
In Fig. 8 (c), in contrast to two-stage 3D VG methods, one-stage 3D VG methods integrate object detection and feature extraction guided by language queries, making it easier to locate objects.
3D-SPS [141] took the 3D VG task as a keypoint selection problem and avoided the separation of detection and matching. Specifically, 3D-SPS initially coarsely samples language-related keypoints through the description-aware keypoint sampling module. Subsequently, it finely selects target keypoints and predicts the foundation using the goal-oriented progressive mining module. Inspired by the 2D image language pre-train model such as MDETR[155] and GLIP[156], BUTD-DETR[142] proposed a bottom-up top-down detection transformer that can be used for 2D and 3D VG. Concretely, BUTD-DETR utilizes labeled bottom-up box proposals and top-down language descriptions to guide the decoding of target objects and corresponding language spans through the prediction head.
However, these methods either extract sentence-level features that couple all words or focus more on object names in the description, which would lose the word-level information or neglect other attributes. To address these issues, EDA[143] explicitly decoupled the textual attributes in a sentence and conducted dense alignment between fine-grained language and point cloud objects. Firstly, the long text is decoupled into five semantic components, including main object, auxiliary object, attributes, pronoun, and relation. Subsequently, the dense alignment is designed to align all object-related decoupled textual semantic components with visual features. To reason human intentions from implicit instructions, ReGround3D [144] designed a visual-centric reasoning module, powered by a MLMs, and a 3D grounding module that accurately obtains object locations by revisiting enhanced geometry and fine-grained details from 3D scenes. Additionally, a Chain-of-Grounding mechanism is employed to improve 3D reasoning grounding through interleaved reasoning and grounding steps.
IV-C Visual Language Navigation
@@
Visual Language Navigation (VLN) stands as a key research problem of Embodied AI, aiming at enabling agents to navigate in unseen environments following linguistic instructions. VLN requires robots to understand complex and diverse visual observations and meanwhile interpret instructions at different granularities. The input for VLN typically consists of two parts: visual information and natural language instructions. The visual information can either be a video of past trajectories or a set of historical-current observation images. The natural language instructions include the target that the embodied agent needs to reach or the task that the embodied agent is expected to complete. The embodied agent must use the above information to select one or a series of actions from a list of candidates to fulfill the requirements of the natural language instructions. This process could be represented as:
$Action=\mathcal{M}(O,H,I)$ (1)
where $Action$ is the chosen action or a list of action candidates, $O$ is the current observation, $H$ is the historical information, and $I$ is the natural language instruction.
TABLE VI: Comparison of Different VLN Datasets.
@@
R2R [63] 2018 Matterport3D Indoor, Discrete Step-by-step instructions 21,567
R4R [157] 2019 Matterport3D Indoor, Discrete Step-by-step instructions 200,000+
VLN-CE [158] 2020 Habitat Indoor, Continuous Step-by-step instructions -
TOUCHDOWN [159] 2019 - Outdoor, Discrete Step-by-step instructions 9,326
REVERIE [160] 2020 Matterport3D Indoor, Discrete Described goal navigation 21,702
SOON [161] 2021 Matterport3D Indoor, Discrete Described goal navigation 3,848
DDN [162] 2023 AI2-THOR Indoor, Continuous Demand-driven navigation 30,000+
ALFRED [163] 2020 AI2-THOR Indoor, Continuous Navigation with interaction 25,743
OVMM [164] 2023 Habitat Indoor, Continuous Navigation with interaction 7,892
BEHAVIOR-1K[165] 2023 OmniGibson Indoor, Continuous Long-span navigation with interaction, 1,000
CVDN [166] 2020 Matterport3D Indoor, Discrete Dialog, oracle 2,050
DialFRED [167] 2022 AI2-THOR Indoor, Continuous Dialog, oracle 53,000
SR (Success Rate), TL (Trajectory Length), and SPL (Success Weighted by Path Length) are the most commonly used metrics in VLN. Among them, SR directly reflects the navigation performance of the embodied agent, TL reflects the navigation efficiency, and SPL combines both to indicate the overall performance of the embodied agent. Below, we introduce VLN in two parts: datasets and methods.
IV-C1 Datasets
@@
Room to Room (R2R) [63] is a VLN dataset based on Matterport3D. In R2R, embodied agents navigate according to step-by-step instructions, choosing the next adjacent navigation graph node to advance based on visual observations until they reach the target location. Embodied agents need to dynamically track progress to align the navigation process with fine-grained instructions. Room-for-Room [157] extends the paths in R2R to longer trajectories, which requires stronger long-distance instruction and history alignment capabilities of embodied agents. VLN-CE [158] extends R2R and R4R to continuous environments, embodied agents can move freely in the scene. This makes the action decision of embodied agents more difficult. Different from the above datasets based on indoor scenes, TOUCHDOWN dataset [159] is created based on Google Street View. In TOUCHDOWN, embodied agents follow instructions to navigate in the street view rendering simulation of New York City to find the specified object.
@@
Similar to R2R, the REVERIE dataset [160] is also built based on the Matterport3D simulator. REVERIE requires embodied agents to accurately locate the distant invisible target object specified by concise, human-annotated high-level natural language instructions, which means that embodied agents need to find the target object among a large number of objects in the scene. In SOON [161], agents receive a long and complex instruction from coarse to fine to find the target object in the 3D environment. During navigation, agents first search a larger area, and then gradually narrow the search range according to the visual scene and instructions. This makes SOON’s navigation target-oriented and independent of the initial position. DDN [162] moves a step further beyond these datasets, only providing human demands without specifying explicit objects. The agent needs to navigate through the scene to find objects that meet human needs.
@@
ALFRED dataset [163] is based on the AI2-THOR simulator. In ALFRED, embodied agents need to understand environmental observations and complete household tasks in an interactive environment according to coarse-grained and fine-grained instructions. The task of OVMM [164] is to pick any object in any unseen environment and place it in a specified location. Agents need to locate the target object in the home environment, navigate and grab it, and then navigate to the target location to put down the object. OVMM provides a simulation based on Habitat and a framework for implementation in the real world. Behavior-1K dataset [165] is based on human needs, comprising 1,000 long-sequence, complex, skill-dependent daily tasks, which are designed within OmniGibson, an extension of the iGibson simulation environment. Agents need to complete long-span navigation-interaction tasks which contain thousands of low-level action steps based on visual information and language instructions. These complex tasks requires strong capabilities of understanding and memory.
There are also some more special datasets. CVDN [166] requires embodied agents to navigate to the target based on dialogue history, and ask questions for help to decide the next action when uncertain. DialFRED [167], an extension of ALFRED, allows agents to ask questions during the navigation and interaction process to get help. These datasets all introduce additional oracles, and embodied agents need to obtain more information beneficial to navigation by asking questions.
IV-C2 Method
@@
TABLE VII: Comparison of VLN methods.
@@
Memory-Understanding based. Graph-based learning is an essential part of the memory-understanding based method. Graph-based learning usually represents the navigation process in the form of a graph, where the information obtained by the embodied agent at each time step is encoded as nodes of the graph. Embodied agent obtains global or partial navigation graph information as a representation of the historical trajectory. LVERG [168] encoded the language information and visual information of each node separately, design a new language and visual entity relationship graph to model the inter-modal relationship between text and vision, and the intra-modal relationship between visual entities. LM-Nav [172] used a goal-conditioned distance function to infer connections between original observation sets and construct a navigation graph, and extracts landmarks from the instructions through a LLM, uses a visual language model to match them with the nodes of the navigation graph. Although HOP [173] is not based on graph learning, its method is similar to the graph, requiring the model to model time-ordered information at different granularities, thereby achieving a deep understanding of historical trajectories and memories.
@@
Different learning schemes explore how to utilize historical trajectories and memories better. Through adversarial learning, CMG [169] alternated between imitation learning and exploration encouragement schemes, effectively strengthening the understanding of instructions and historical trajectories, shortening the difference between training and inference. GOAT [177] directly trained unbiased models through Backdoor Adjustment Causal Learning (BACL) and Frontdoor Adjustment Causal Learning (FACL), conducts contrastive learning with vision, navigation history, and their combination to instructions, enabling the agent to make fuller use of information. The enhanced cross-modal matching method proposed by RCM [170] used goal-oriented external rewards and instruction-oriented internal rewards to perform cross-modal grounding globally and locally and learns from its own historical good decisions through self-supervised imitation learning. FSTT [175] introduced TTA into VLN and optimizes the model in terms of gradients and model parameters at two scales of time steps and tasks, effectively improving model performance.
The specific application of large models in Memory-Understanding based methods is to understand the representation of historical memory and to understand the environment and tasks based on its extensive world knowledge. NaviLLM [174] integrated the historical observation sequence into the embedding space through the visual encoder, inputs the multi-modal information of the fusion encoding into the LLM and fine-tunes it, reaching the state-of-the-art on multiple benchmarks. NaVid [179] maked improvements in the encoding of historical information, achieves different degrees of information retention on historical observations and current observations through different degrees of pooling. DiscussNav [176] assigned large model experts with different abilities to different roles, drives the large models to discuss before navigation actions to complete navigation decisions, and achieves excellent performance in zero-shot VLN.
@@
IV-D Non-Visual Perception: Tactile
Tactile sensors provide agents with detailed information such as texture, hardness and temperature. For the same action, the knowledge learned from vision and tactile sensors may be related and complementary, allowing robots to fully grasp the high-precision tasks in hand. Tactile perception is therefore vital for agents in the physical world and undoubtedly enhances human-computer interaction[188, 189, 190].
For tactile perception tasks, the agent needs to gather tactile information from the physical world and then perform complex tasks. In this section, as shown in Fig. 10, we first introduce the existing types of tactile sensors and their datasets, and then discuss three main tasks in tactile perception: estimation, recognition, and manipulation.
Figure 10: Different types of Tactile Sensors. Non-vision sensors (a) mainly use sensors of force, pressure, vibration and temperature to get tactile knowledge. Vision-based tactile sensors ((b)-(e)) are based on optical principles. A camera is placed behind the gel to record the image of its deformation, using illumination from light sources at different directions. (a)-(e) are the details from BioTac, Gelsight, DIGIT, 9DTact and Gelsilm.
IV-D1 Sensor Design
The principle of human tactile is that the skin changes shape when touched, and its abundant nerve cells send electrical signals, which also serves as the basis for designing tactile sensors. Tactile sensor design methods can be divided into three categories: non-vision-based, vision-based, and multi-modal. Non-vision-based tactile sensors, primarily using electrical and mechanical principles, chiefly register fundamental, low-dimensional sensory outputs such as force, pressure, vibration, and temperature[191, 192, 193, 194, 195, 196]. One of the notable representatives is BioTac [197] and its simulator[198]. Vision-based tactile sensors are based on optical principles. Using images of the gel’s deformation as tactile information, vision-based tactile sensors such as GelSight[199], Gelslim[200], DIGIT[201], 9DTact[202], TacTip[203], GelTip[204] and AllSight[205] have been used for numerous applications. Simulations like TACTO[206] and Taxim[207] are also popular. Recent work has focused on cost reduction[202] and integration into robotic hands[201, 208, 209]. Multi-modal tactile sensors, inspired by human skin, combine multi-modal information like pressure, proximity, acceleration, and temperature, using flexible materials and modular design.
IV-D2 Datasets
The datasets of non-vision sensors, mainly collected by BioTac series[197], contain electrode values, force vectors and contact location. Since tasks are mainly the estimation of force and grasping details, objects in datasets are usually force and grasping samples. Vision-based sensors, with high-resolution images of deformation gel, focus more on higher estimation, texture recognition and manipulation. The datasets are collected by Geisight sensors, DIGIT sensors and their simulators[199, 201, 206, 202], consisting household object, wildlife environments, different materials and grasping items. Since image information can be easily aligned and bound with the other modalities (images, language, audio, etc)[14, 210], tactile perception in embodied agent mainly revolves around visual-based sensors. We introduce ten main tactile datasets, which are summarized in Table VIII.
TABLE VIII: Comparison of different tactile datasets.
Dataset Year Sensor Data Format Type Real/ Sensor/ Size Continuous
Types Simulated Simulator
The Feeling of Success[211] 2017 Vision Tac, Vision, Label Daily Necessities Real Gelsight 106 Objects $\surd$
SynTouch[198] 2019 Non-vision Tac, Force, Location - Real BioTac 300k Readings -
Decoding the BioTac[212] 2020 Non-vision Tac, Force, Location - Real BioTac 50k Points $\surd$
ObjectFolder 1.0[213] 2021 Vision Tac, Vision, Audio Household Object Simulate TACTO 100 Objects $\surd$
ObjectFolder 2.0[214] 2022 Vision Tac, Vision, Audio Household Object Simulate Taxim 1000 Objects $\surd$
[…]
YCB-Slide[216] 2022 Vision Tac, Vision, Other Daily Necessities Both TACTO&DIGIT 10 Objects $\surd$
ObjectFolder Real[217] 2023 Vision Tac, Vision, Audio Household Objects Real Gelslim 100 Objects $\surd$
[…]
TaRF[219] 2024 Vision Tac, Vision Full Scene Simulate DIGIT 19.3k -
IV-D3 Methods
Tactile perception has numerous applications and can be categorized into three types: estimation, precise robotic manipulation, and multi-modal recognition tasks.
Estimation
Early works in estimation predominantly focused on basic algorithm for shape, force and slip measurement [220, 221, 202]. Researchers simply used a threshold or applied Convolutional Neural Networks (CNN) to settle these task, based on the colors of tactile images and the change of markers distributions at different frames. The emphasis of the estimation work is mainly on the second stage, the generation of tactile image and the reconstruction of the object. The generation of tactile image [222, 223, 224, 225] aims at generating tactile image from vision data. At first it applied deep learning model that takes RGB-D images as input and outputted tactile images [222, 223]. Recently, as the image generation developed fast, Higuera et al. [224] and Yang et al. [225] applied diffusion model on tactile generation and it performed well. The reconstruction of the object can be divided into 2D reconstruction [226, 227] and 3D reconstruction [228, 229, 202, 230, 231, 232, 233, 234, 235, 236, 219, 237, 238, 239, 240, 241]. 2D reconstruction mainly focuses on the shape and segmentation of the object while 3D focuses on the surface and pose, and then even the full scene perception. The tasks first adopted mathematics methods, auto-encoder methods and neural network methods to fuse visual (sometime point clouds) and tactile features together. Lately, researchers like Comi et al. [236] and Dou et al. [219] implemented new methods based on Neural Radiance Fields (NeRF) and 3D Gaussian Splatting (3DGS) into tactile reconstruction work.
Robotic Manipulation
In tactile task, bridging the sim-to-real gap is more than important. Reinforcement learning and GAN-based methods have been proposed to address variations in accurate, on-time robotic manipulation tasks.
Reinforcement Learning method. Visuotactile-RL[242] proposed several methods to existing RL methods, including tactile gating, tactile data augmentation and visual degradation. Rotateit[243] is a system that enables fingertip-based object rotation along multiple axes by leveraging multimodal sensory inputs. It trained the network by reinforcement learning policies with privileged information and enabled online inference. [244] proposed a deep RL approach to object pushing using only tactile perception. It came up with a goalconditioned formulation that allows both model-free and modelbased RL to obtain accurate policies for pushing an object to a goal. AnyRotate[245] focused on in-hand manipulation. It is a system for gravity-invariant multi-axis in-hand object rotation using dense featured sim-to-real touch, constructing a continuous contact feature representation to provide tactile feedback for training a policy in simulation and introduce an approach to perform zero-shot policy transfer by training an observation model to bridge the sim-to-real gap.
GAN-based method. ACTNet[246] proposed an unsupervised adversarial domain adaptation method to narrow the domain gap for pixel-level tactile perception tasks. An adaptively correlative attention mechanism was introduced to improve the generator, which is capable of leveraging global information and focusing on salient regions. However, pixel-level domain adaptation lead to error accumulation, degrade performance, and increased structural complexity and training costs. In comparison, STR-Net[247] proposed a feature-level unsupervised framework for tactile images, narrowing the domain gap for feature-level tactile perception tasks. Moreover, some methods focused on sim-to-real. For example, the Tactile Gym 2.0 [248]. However, due to its complexity and high cost, it is challenging for practical application.
Recognition
Tactile representations learning focuses on material classification and multi-modal understanding, which can be divided into two categories: Traditional Methods and LLMs&VLMs Methods.
Traditional Methods. Various traditional approaches have been employed to enhance tactile representation learning. Autoencoder frameworks have been instrumental in developing compact tactile data representations. Polic et al.[249] used a convolutional neural network autoencoder for dimensionality reduction of optical-based tactile sensor images. Gao et al. [250] created a supervised recurrent autoencoder to handle heterogeneous sensor datasets, while Cao et al.[251] created TacMAE used a masked autoencoder for incomplete tactile data. Zhang et al.[252] introduced MAE4GM, a multimodal autoencoder integrating visuo-tactile data. Since tactile acts as a complement to other modes, Joint Training methods are used to fuse multiple modalities. Yuan et al.[253] trained CNNs with modalities that included depth, vision, and tactile data. Similarly, Lee et al. [254] used a variational Bayesian approach for modalities like force sensors series and end-effector metrics. For better learning representation, Self-supervised methods like contrastive learning are also key techniques in binding modalities together. Researches differ in contrastive methods Lin et al. [255] simply paired tactile inputs with multiple visual inputs and Yang et al. [256] employed visuo-tactile contrastive multiview features. Kerr et al. [215] used InfoNCE loss and Guzey et al. [257] used BYOL. These traditional methods have established a solid foundation for tactile representation learning.
LLMs&VLMs Methods. LLM and VLM have shown an amazing understanding of cross-modal interactions and strong zero-shot performance recently. Recent works from Yang et al.[189], Fu et al. [218] and Yu et al. [258] encoded and aligned tactile data with visual and language modalities by contrastive pretrained method. Then an LLM like LLaMA would be applied, using a fine-tune method to fit tasks like tactile description. The advent of LLM and VLM techniques has further advanced the field, enabling more comprehensive and robust cross-modal tactile representations.
IV-D4 Difficulties
a) Disadvantages of different sensor types: Traditional sensors deliver simple and low-dimensional data, challenging multi-modal learning. Vision-based sensors and electronic skins, while highly accurate, are expensive. b) Data acquisition challenges: Collecting data, especially both tactile and visual simultaneously, is difficult despite some progress in developing simplified collection devices. c) Inconsistent standards: Tactile sensors operate with inconsistent standards and principles, hindering large-scale learning and limiting the usefulness of public datasets. There is a need for standardized and extensive datasets.

### v8

@@ IV-B Visual Language Navigation
Visual Language Navigation (VLN) is an essential task, aiming at navigating in unseen environments following linguistic instructions. VLN requires robots to understand complex and diverse visual observations and meanwhile interpret instructions at different granularities. The input typically consists of two parts: visual information and natural language instructions. The visual information can either be a video of past trajectories or a set of historical-current observation images. The natural language instructions include the target that the agent needs to reach or the task that the agent is expected to complete. The agent must use the above information to select one or a series of actions from a list of candidates to fulfill the requirements of the natural language instructions. This process could be represented as $Action=\mathcal{M}(O,H,I)$, where $Action$ is the chosen action or a list of action candidates, $O$ is the current observation, $H$ is the historical information, and $I$ is the natural language instruction.
TABLE IV: Comparison of Different VLN Datasets. M3D: Matterport3D, AT: AI2-THOR, OG: OmniGibson, I: Indoor, D: Discrete, O: Outdoor, C: Continuous, SbS: Step-by-step instructions, DGN: Described goal navigation, DDN: Demand-driven navigation, NwI: Navigation with interaction, LSNwI: Long-span navigation with interaction, D&O: Dialog and oracle
@@
R2R [105] 2018 M3D I, D SbS 21,567
R4R [106] 2019 M3D I, D SbS 200,000+
VLN-CE [107] 2020 Habitat I, C SbS -
TOUCHDOWN [108] 2019 - O, D SbS 9,326
REVERIE [109] 2020 M3D I, D DGN 21,702
SOON [110] 2021 M3D I, D DGN 3,848
DDN [111] 2023 AT I, C DDN 30,000+
ALFRED [112] 2020 AT I, C NwI 25,743
OVMM [113] 2023 Habitat I, C NwI 7,892
BEHAVIOR-1K[114] 2023 OG I, C LSNwI 1,000
CVDN [115] 2020 M3D I, D D&O 2,050
DialFRED [116] 2022 AT I, C D&O 53,000
IV-B1 Datasets
@@
Room to Room (R2R)[105] is a VLN dataset based on Matterport3D. In R2R, embodied agents navigate according to step-by-step instructions, choosing the next adjacent navigation graph node to advance based on visual observations until they reach the target location. Room-for-Room [106] extends the paths in R2R to longer trajectories, which requires stronger long-distance instruction and history alignment capabilities of embodied agents. VLN-CE [107] extends R2R and R4R to continuous environments, embodied agents can move freely in the scene. Different from the above datasets based on indoor scenes, the TOUCHDOWN dataset [108] is created based on Google Street View. In TOUCHDOWN, embodied agents follow instructions to navigate in the street view rendering simulation of New York City to find the specified object. Similar to R2R, the REVERIE dataset [109] is also built based on the Matterport3D simulator. REVERIE requires embodied agents to accurately locate the distant invisible target object specified by concise, human-annotated high-level natural language instructions. In SOON [110], agents receive a long and complex instruction from coarse to fine to find the target object in the 3D environment. During navigation, agents first search a larger area, and then gradually narrow the search range according to the visual scene and instructions. DDN [111] moves a step further beyond these datasets, only providing human demands without specifying explicit objects. The agent needs to navigate through the scene to find objects.
@@
ALFRED dataset [112] is based on the AI2-THOR simulator. In ALFRED, embodied agents need to understand environmental observations and complete household tasks in an interactive environment according to coarse-grained and fine-grained instructions. The task in OVMM [113] dataset is to pick any object in any unseen environment and place it in a specified location. OVMM provides a simulation based on Habitat and a framework for implementation in the real world. Behavior-1K dataset [114] is based on human needs, comprising 1,000 long-sequence, complex, skill-dependent daily tasks. Agents need to complete long-span navigation-interaction tasks which contain thousands of low-level action steps based on visual information and language instructions. These complex tasks requires strong capabilities of understanding and memory. CVDN [115] requires embodied agents to navigate to the target based on dialogue history, and ask questions for help to decide the next action when uncertain. DialFRED [116], an extension of ALFRED, allows agents to ask questions during the navigation and interaction process to get help. These datasets introduce additional oracles, and embodied agents need to obtain more information beneficial to navigation by asking questions.
IV-B2 Method
@@
TABLE V: Comparison of VLN methods.
@@
Memory-Understanding based. Graph-based learning is an essential part of the memory-understanding based method. It usually represents the navigation process in the form of a graph, where the information obtained by the agent at each time step is encoded as nodes of the graph. The agent obtains global or partial navigation graph information as a representation of the historical trajectory. LVERG [117] encoded the language information and visual information of each node separately, design a new language and visual entity relationship graph to model the inter-modal relationship between text and vision, and the intra-modal relationship between visual entities. LM-Nav [121] used a goal-conditioned distance function to infer connections between original observation sets and construct a navigation graph, and extracted landmarks from the instructions through a LLM. Although HOP [122] is not based on graph learning, it requires to model time-ordered information at different granularities, thereby achieving a deep understanding of historical trajectories and memories.
@@
Different learning schemes explore how to utilize historical trajectories and memories better. Through adversarial learning, CMG [118] alternated between imitation learning and exploration encouragement schemes, effectively strengthening the understanding of instructions and historical trajectories, shortening the difference between training and inference. GOAT [126] directly trained unbiased models through Backdoor Adjustment Causal Learning (BACL) and Frontdoor Adjustment Causal Learning (FACL), conducted contrastive learning with vision, navigation history, and their combination to instructions, enabling the agent to make fuller use of information. The enhanced cross-modal matching method proposed by RCM [119] used goal-oriented external rewards and instruction-oriented internal rewards to perform cross-modal grounding globally and locally and learns from its own historical good decisions through self-supervised imitation learning. FSTT [124] introduced TTA into VLN and optimizes the model in terms of gradients and model parameters at two scales of time steps and tasks, effectively improving model performance.
The specific application of large models in Memory-Understanding based methods is to understand the representation of historical memory and to understand the environment and tasks based on its extensive world knowledge. NaviLLM [123] integrated the historical observation sequence into the embedding space through the visual encoder, inputs the multi-modal information of the fusion encoding into the LLM and fine-tunes it, reaching the state-of-the-art on multiple benchmarks. NaVid [128] made improvements in the encoding of historical information, achieves different degrees of information retention on historical observations and current observations through different degrees of pooling. LH-VLN [137] proposed the NavGen platform, the long-horizon navigation benchmark, and the Multi-Granularity Dynamic Memory (MGDM) module to enhance task evaluation and model adaptability in dynamic environments.